# Trivial Isomorphisms between Reduced Products

### v2

@@ 1. Introduction
Remarkably, S. Shelah has proven that in some forcing extension of the universe the situation is totally different: all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial ([33]). An automorphism $\Phi$ of ${\mathcal{P}}({\mathbb{N}})/{\mathcal{I}}$ is trivial if there are sets $A$ and $B$ whose complements are in ${\mathcal{I}}$ and a bijection $f\colon A\to B$ such that $X\mapsto f[X]$ lifts $\Phi$. It is not difficult to see that this is equivalent to asserting that $\Phi$ has a lifting that is a continuous homomorphism. Shelah’s conclusion has since been deduced from the Proper Forcing Axiom ([35]) and from the conjunction of two of its consequences $\operatorname{{\mathsf{OCA}_{T}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$ ([39]). While $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$ alone does not suffice for this conclusion by [39, Theorem 3.1], until now it was open whether $\operatorname{{\mathsf{OCA}_{T}}}$ alone does.
For general reduced products and other quotient structures, isolating the correct notion of a trivial automorphism (or trivial isomorphism) is a nontrivial issue. For example, in case of the Calkin algebra, trivial automorphisms coincide with inner automorphisms ([11], see also [12, §17]), but for other corona $\mathrm{C^{*}}$-algebras many trivial automorphisms are outer ([40]). In case of Čech–Stone remainders, the situation is analogous to that of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ ([40, Theorem C]). In the present paper we will not delve into $\mathrm{C^{*}}$-algebras or other metric structures.
@@
Towards a strong version of rigidity for isomorphisms of reduced products, we introduce a sharpening of Todorčević’s Open Colouring Axiom, $\operatorname{{\mathsf{OCA}_{T}}}$ (Definition 3.2). This axiom, denoted $\operatorname{{\mathsf{OCA}^{\#}}}$, substantially simplifies the ‘usual’ uniformization arguments needed when proving rigidity results (see e.g. [8, forcings ${\mathcal{P}}$ and ${\mathcal{P}}_{\omega_{1}}$ in §3], [29], or [13]). We prove that $\operatorname{{\mathsf{OCA}^{\#}}}$ alone implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial, and that it suffices even for the case ${\mathcal{I}}=\operatorname{Fin}$ of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem of [8], originally proven by an involved uniformization argument using Martin’s Axiom (see Theorem 4.7). Even better, we adapt the proof in [30, §5] to show that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$, therefore giving the following strengthening of the main result of [39]:
@@
$\operatorname{{\mathsf{OCA}_{T}}}$ implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial.
@@
Assume $\operatorname{{\mathsf{OCA}_{T}}}+\operatorname{{\mathsf{MA}_{\aleph_{1}}(\sigma\textrm{-linked})}}$. Then every coordinate respecting isomorphism between reduced products modulo $\operatorname{Fin}$ of structures in the same countable language is trivial.
@@
In §2 we introduce the notions of coordinate respecting functions and of theories which recognize coordinates. A reduced product $\prod_{n\in{\mathbb{N}}}M_{n}/{\mathcal{I}}$ of models of a theory that recognizes coordinates contains a definable copy of ${\mathcal{P}}({\mathbb{N}})/{\mathcal{I}}$, or at least a workable poor man’s version of one. [^7] The analogous assertion applies to reduced products over larger index sets, but we will not explore consequences of this observation in the present paper. An isomorphism between such reduced products gives an isomorphism between these definable Boolean algebras, and is therefore coordinate respecting. [^8] This is not nearly as straightforward as it may sound; see Definition 2.25. We then prove that theories of fields, linear orders, trees, and certain graphs recognize coordinates. Sections 3 and 4 are dedicated to the axiom $\operatorname{{\mathsf{OCA}^{\#}}}$: after introducing it, we prove that it follows from $\operatorname{{\mathsf{OCA}_{T}}}$ and that it implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial. This provides a proof of Theorem 1, and even that $\operatorname{{\mathsf{OCA}_{T}}}$ implies the conclusion of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem ([8, Theorem 3.3.5]) for $\operatorname{Fin}$. Section 5 is dedicated to the proof of our main technical result, Theorem 7, and in §6 we prove Theorems 25. In §6.2 we analyse the (lack of) rigidity under CH; these results are straightforward and included only for completeness. Lastly, §7 is an intriguing footnote to our main results. In it we give situations in which automorphisms (and isomorphisms) of arbitrary (not reduced!) products are products of isomorphisms of the fibers. These results rely on a unique factorization result for products and are, to the best of our knowledge, new. Appendix A contains the proof of a well-known, but perhaps not so well-documented, descriptive set theoretic fact used in our uniformization arguments in §4 and §5.
@@ 2. Coordinate respecting functions
The conclusion of Lemma 2.2 (3) can fail if the finiteness assumption is dropped. More precisely, there are a language $\mathcal{L}$, $\mathcal{L}$-structures $\mathcal{M}_{n}$, and a trivial automorphism of product form $\Phi\colon\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\to\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}$ such that it is not possible to choose all but finitely many of $h_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{M}}_{n}$ to be homomorphisms. Let $\mathcal{L}$ be the language with infinitely many unary predicates $P_{n}(x)$, for $n\in{\mathbb{N}}$. For each $n$ take ${\mathcal{M}}_{n}={\mathcal{N}}_{n}$ to be equal to the unique countable model ${\mathcal{M}}$ in which $P_{k}^{\mathcal{M}}$, for $k\in{\mathbb{N}}$, form an independent family (i.e. the set $\bigcap_{k\in K}P_{k}^{\mathcal{M}}\cap\bigcap_{l\in L}M\setminus P_{l}^{\mathcal{M}}$ is nonempty for every two disjoint finite subsets $K,L$ of ${\mathbb{N}}$). For a fixed $n$ the sets
@@ 2.2.2. Graphs
(2.4) $(\forall^{{\mathcal{I}}}n)\,E^{{\mathcal{M}}_{n}}(\tilde{a}_{n},\tilde{b}_{n}).$
@@
(2.5) $(\forall^{\mathcal{I}}n)\,E^{{\mathcal{M}}_{n}}(\tilde{a}_{n},\tilde{b}_{n})\wedge\tilde{a}_{n}\neq\tilde{b}_{n}.$
@@
(2.7) $a\perp b\text{ if }\pi_{S}(a)\neq\pi_{S}(b)\text{ for all nonzero }S\in{\mathcal{P}}({\mathbb{N}})/{\mathcal{I}}.$
@@
If each ${\mathcal{M}}_{n}$ is a graph with at least two vertices and ${\mathcal{M}}_{n}$ is the null graph for infinitely many $n$, then for any two distinct vertices $a$ and $b$ in ${\mathcal{M}}=\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}$ (2.5) holds and therefore ${\mathcal{M}}$ is also a null graph. In this case every definable set in ${\mathcal{M}}$ is finite or cofinite. Therefore neither the sets $A_{ab}$ nor the relation $\perp$ are definable and ${\mathcal{M}}$ has $2^{2^{\aleph_{0}}}$ automorphisms in $\operatorname{{\mathsf{ZFC}}}$.
The random graph $\mathcal{R}$ (also known as the Rado graph) is the unique graph with countable set of vertices such that for every two disjoint finite sets of vertices $P$ and $Q$ there is a vertex $v$ adjacent to all vertices in $P$ and not adjacent to the unique vertex in $Q$.
@@
To prove the direct implication in the first assertion, assume $x\in A_{ab}$. Then $(\forall^{\mathcal{I}}n)\,\tilde{x}_{n}\in\{\tilde{a}_{n},\tilde{b}_{n}\}$. If $z\in{\mathcal{M}}$ satisfies $E(z,a)$ and $E(z,b)$, then the characterization of adjacency given in (2.4) implies $E(z,x)$.
For the converse implication, assume $x\notin A_{ab}$. Then $S=\{n\mid\tilde{x}_{n}\notin\{\tilde{a}_{n},\tilde{b}_{n}\}\}$ is ${\mathcal{I}}$-positive. Since each ${\mathcal{M}}_{n}$ is sufficiently random, for each $n\in S$ there are $v_{n}$ in ${\mathcal{M}}_{n}\setminus\{\tilde{a}_{n},\tilde{b}_{n}\}$ and $\tilde{z}_{n}\in{\mathcal{M}}_{n}$ such that $\lnot E(\tilde{z}_{n},v_{n})$, $E(\tilde{z}_{n},\tilde{a}_{n})$, and $E(\tilde{z}_{n},\tilde{b}_{n})$. This implies $\lnot E(z,x)$, hence the right-hand side of the first assertion fails.
@@
Vice versa, suppose that $a\not\perp b$, and let $\tilde{S}=\{n\mid\tilde{a}(n)=\tilde{b}(n)\}$. Since $a$ and $b$ are distinct, the complement of $\tilde{S}$ is not in ${\mathcal{I}}$. Let $\tilde{c}$ be such that $\tilde{c}(n)\neq\tilde{a}(n)$ for all $n\in\tilde{S}$ and $\tilde{c}(n)=\tilde{b}(n)$ for $n\notin\tilde{S}$. Then $b\in A_{ac}$, but clearly $b\neq c$. ∎
@@
Next, fix $a,b$ and $c$ in ${\mathcal{M}}$ with $(a,b)$ and $(a,c)$ in ${\mathcal{Z}}$. We claim that $\alpha^{ab}=\alpha^{ac}$. Assume otherwise and fix $S$ such that $\alpha^{ab}(S)\cap\alpha^{ac}(S)=\emptyset$. Lift $S$ to $\tilde{S}\in{\mathcal{P}}({\mathbb{N}})$. Let $d$ be constructed in such a way that:
@@
Such a $d$ exists as each graph is sufficiently random. Note that $\neg E(a,d)$. On the other hand, we have $E(d,\chi_{S}^{ab})$ and $E(d,\chi_{S}^{ac})$, henceforth $E(\Phi(d),\chi_{\alpha^{ab}(S)}^{\Phi(a)\Phi(b)})$ and $E(\Phi(d),\chi_{\alpha^{ac}(S)}^{\Phi(a)\Phi(c)})$. As $\alpha^{ab}(S)$ and $\alpha^{ac}(S)$ are disjoint, the union of their complements is (modulo ${\mathcal{J}}$) equal to $\mathbb{N}$. Lifting $\Phi(d),\chi_{\alpha^{ac}(S)}^{\Phi(a)\Phi(c)},\chi_{\alpha^{ab}(S)}^{\Phi(a)\Phi(b)}$ respectively, and using the fact that $E(x,y)$ if and only if $\forall^{\mathcal{J}}n\,(E(\tilde{x}(n),\tilde{y}(n)))$, we have that $E(\Phi(a),\Phi(d))$, which is a contradiction.
@@ 2.3. A general local criterion
A theory $T$ recognizes coordinates locally if there is a set of pairs ${\mathcal{Z}}$ definable in $T^{*}$ such that for all $(a,b)\in{\mathcal{Z}}$, we have that $a\perp b$ and the sets $A_{ab}$ (as in (2.6)), as well as their canonical Boolean algebra structure, are uniformly definable [^13] Uniformly definable here means that the same formulas define $A_{ab}$, as well as its canonical Boolean algebra structure, for all relevant pairs $a,b$ in all relevant reduced products. in all reduced products of models of $T$ over atomless ideals.
@@
Let $T$ be a theory that recognizes coordinates locally. Let ${\mathcal{M}}_{n}$ and ${\mathcal{N}}_{n}$ be models of $T$, and let ${\mathcal{I}}$ and ${\mathcal{J}}$ be atomless ideals on ${\mathbb{N}}$. Let ${\mathcal{M}}=\prod_{n}{\mathcal{M}}_{n}/{\mathcal{I}}$ and ${\mathcal{N}}=\prod_{n}{\mathcal{N}}_{n}/{\mathcal{J}}$, and suppose that $\Phi\colon{\mathcal{M}}\to{\mathcal{N}}$ is an isomorphism. Fix ${\mathcal{Z}}$ as in Definition 2.25.

### v3

@@ 1. Introduction
Remarkably, S. Shelah has proven that in some forcing extension of the universe the situation is totally different: all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial ([33]). An automorphism $\Phi$ of ${\mathcal{P}}({\mathbb{N}})/{\mathcal{I}}$ is trivial if there are sets $A$ and $B$ whose complements are in ${\mathcal{I}}$ and a bijection $f\colon A\to B$ such that $X\mapsto f[X]$ lifts $\Phi$. It is not difficult to see that this is equivalent to asserting that $\Phi$ has a lifting that is a continuous homomorphism. Shelah’s conclusion has since been deduced from the Proper Forcing Axiom ([35]) and from the conjunction of two of its consequences $\operatorname{{\mathsf{OCA}_{T}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$ ([38]). While $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$ alone does not suffice for this conclusion by [38, Theorem 3.1], until now it was open whether $\operatorname{{\mathsf{OCA}_{T}}}$ alone does.
For general reduced products and other quotient structures, isolating the correct notion of a trivial automorphism (or trivial isomorphism) is a nontrivial issue. For example, in case of the Calkin algebra, trivial automorphisms coincide with inner automorphisms ([11], see also [12, §17]), but for other corona $\mathrm{C^{*}}$-algebras many trivial automorphisms are outer ([39]). In case of Čech–Stone remainders, the situation is analogous to that of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ ([39, Theorem C]). In the present paper we will not delve into $\mathrm{C^{*}}$-algebras or other metric structures.
@@
Towards a strong version of rigidity for isomorphisms of reduced products, we introduce a sharpening of Todorčević’s Open Colouring Axiom, $\operatorname{{\mathsf{OCA}_{T}}}$ (Definition 3.2). This axiom, denoted $\operatorname{{\mathsf{OCA}^{\#}}}$, substantially simplifies the ‘usual’ uniformization arguments needed when proving rigidity results (see e.g. [8, forcings ${\mathcal{P}}$ and ${\mathcal{P}}_{\omega_{1}}$ in §3], [29], or [13]). We prove that $\operatorname{{\mathsf{OCA}^{\#}}}$ alone implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial, and that it suffices even for the case ${\mathcal{I}}=\operatorname{Fin}$ of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem of [8], originally proven by an involved uniformization argument using Martin’s Axiom (see Theorem 4.7). Even better, we adapt the proof in [30, §5] to show that $\operatorname{{\mathsf{OCA}_{T}}}$ implies $\operatorname{{\mathsf{OCA}^{\#}}}$, therefore giving the following strengthening of the main result of [38] and of of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem ([8, Theorem 3.3.5]) for $\operatorname{Fin}$.
@@
$\operatorname{{\mathsf{OCA}_{T}}}$ implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial. It moreover implies that every homomorphism $\Phi\colon{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}\to{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ has a decomposition $\Phi=\Phi_{1}\oplus\Phi_{2}$, where $\Phi_{1}$ has a completely additive lifting and where the kernel of $\Phi_{2}$ is nonmeager.
@@
Assume $\operatorname{{\mathsf{OCA}_{T}}}+\operatorname{{\mathsf{MA}_{\aleph_{1}}(\sigma\textrm{-linked})}}$. Then every coordinate respecting isomorphism between reduced products modulo $\operatorname{Fin}$ of countable structures in the same language is trivial.
@@
In §2 we introduce the notions of coordinate respecting functions and of theories which recognize coordinates. A reduced product $\prod_{n\in{\mathbb{N}}}M_{n}/{\mathcal{I}}$ of models of a theory that recognizes coordinates contains a definable copy of ${\mathcal{P}}({\mathbb{N}})/{\mathcal{I}}$, or at least a workable poor man’s version of one. [^7] The analogous assertion applies to reduced products over larger index sets, but we will not explore consequences of this observation in the present paper. An isomorphism between such reduced products gives an isomorphism between these definable Boolean algebras, and is therefore coordinate respecting. [^8] This is not nearly as straightforward as it may sound; see Definition 2.25. We then prove that theories of fields, linear orders, trees, and certain graphs recognize coordinates. Sections 3 and 4 are dedicated to the axiom $\operatorname{{\mathsf{OCA}^{\#}}}$: after introducing it, we prove that it follows from $\operatorname{{\mathsf{OCA}_{T}}}$, that it implies all automorphisms of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ are trivial and that it even implies the conclusion of the $\operatorname{{\mathsf{OCA}}}$ lifting theorem ([8, Theorem 3.3.5]) for $\operatorname{Fin}$. This provides a proof of Theorem 1. Section 5 is dedicated to the proof of our main technical result, Theorem 7, and in §6 we prove Theorems 25. In §6.2 we analyse the (lack of) rigidity under CH; these results are straightforward and included only for completeness. Lastly, §7 is an intriguing footnote to our main results. In it we give situations in which automorphisms (and isomorphisms) of arbitrary (not reduced!) products are products of isomorphisms of the fibers. These results rely on a unique factorization result for products and are, to the best of our knowledge, new. Appendix A contains the proof of a well-known, but perhaps not so well-documented, descriptive set theoretic fact used in our uniformization arguments in §4 and §5.
@@ 2. Coordinate respecting functions
The conclusion of Lemma 2.2 (3) can fail if the finiteness assumption is dropped. More precisely, there are a language $\mathcal{L}$, $\mathcal{L}$-structures $\mathcal{M}_{n}$, and a trivial automorphism of product form $\Phi\colon\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}\to\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}$ such that it is not possible to choose all but finitely many of $h_{n}\colon{\mathcal{M}}_{n}\to{\mathcal{M}}_{n}$ to be homomorphisms. Let $\mathcal{L}$ be the language with infinitely many unary predicates $P_{n}(x)$, for $n\in{\mathbb{N}}$. For each $n$ take ${\mathcal{M}}_{n}$ to be equal to the unique countable model ${\mathcal{M}}$ in which $P_{k}^{\mathcal{M}}$, for $k\in{\mathbb{N}}$, form an independent family (i.e. the set $\bigcap_{k\in K}P_{k}^{\mathcal{M}}\cap\bigcap_{l\in L}M\setminus P_{l}^{\mathcal{M}}$ is nonempty for every two disjoint finite subsets $K,L$ of ${\mathbb{N}}$). For a fixed $n$ the sets
@@ 2.2.2. Graphs
(2.4) $(\forall^{{\mathcal{I}}}n)\,E^{{\mathcal{M}}_{n}}(\tilde{a}(n),\tilde{b}(n)).$
@@
(2.5) $(\forall^{\mathcal{I}}n)\,E^{{\mathcal{M}}_{n}}(\tilde{a}(n),\tilde{b}(n))\wedge\tilde{a}(n)\neq\tilde{b}(n).$
@@
(2.7) $a\perp b\text{ if }\pi_{S}(a)\neq\pi_{S}(b)\text{ for all nonzero }S\in{\mathcal{P}}({\mathbb{N}})/{\mathcal{I}},$
@@
If each ${\mathcal{M}}_{n}$ is a graph with at least two vertices and ${\mathcal{M}}_{n}$ is the null graph for infinitely many $n$, then for any two distinct vertices $a$ and $b$ in ${\mathcal{M}}=\prod_{n}{\mathcal{M}}_{n}/\operatorname{Fin}$ (2.5) fails and therefore ${\mathcal{M}}$ is also a null graph. In this case every definable set in ${\mathcal{M}}$ is finite or cofinite. Therefore neither the sets $A_{ab}$ nor the relation $\perp$ are definable and ${\mathcal{M}}$ has $2^{2^{\aleph_{0}}}$ automorphisms in $\operatorname{{\mathsf{ZFC}}}$.
The random graph $\mathcal{R}$ (also known as the Rado graph) is the unique graph with countable set of vertices such that for every two disjoint finite sets of vertices $P$ and $Q$ there is a vertex $v$ adjacent to all vertices in $P$ and not adjacent to any vertex in $Q$.
@@
To prove the direct implication in the first assertion, assume $x\in A_{ab}$. Then $(\forall^{\mathcal{I}}n)\,\tilde{x}(n)\in\{\tilde{a}(n),\tilde{b}(n)\}$. If $z\in{\mathcal{M}}$ satisfies $E(z,a)$ and $E(z,b)$, then the characterization of adjacency given in (2.4) implies $E(z,x)$.
For the converse implication, assume $x\notin A_{ab}$. Then $S=\{n\mid\tilde{x}(n)\notin\{\tilde{a}(n),\tilde{b}(n)\}\}$ is ${\mathcal{I}}$-positive. For each $n\in S$, by applying the definition of sufficiently random to $P=\{\tilde{a}(n),\tilde{b}(n)\}$ and $u=\tilde{x}(n)$, there is $\tilde{z}(n)\in{\mathcal{M}}_{n}\setminus\{\tilde{x}(n)\}$ such that $\lnot E(\tilde{z}(n),\tilde{x}(n))$, $E(\tilde{z}(n),\tilde{a}(n))$, and $E(\tilde{z}(n),\tilde{b}(n))$. Since $S$ is ${\mathcal{I}}$-positive, this implies $\lnot E(z,x)$, hence the right-hand side of the first assertion fails.
@@
Vice versa, suppose that $a\not\perp b$, and let $\tilde{S}=\{n\mid\tilde{a}(n)=\tilde{b}(n)\}$. Since $a$ and $b$ are distinct and $a\not\perp b$, both $\tilde{S}$ and ${\mathbb{N}}\setminus\tilde{S}$ are ${\mathcal{I}}$-positive. Let $\tilde{c}$ be such that $\tilde{c}(n)\neq\tilde{a}(n)$ for all $n\in\tilde{S}$ and $\tilde{c}(n)=\tilde{b}(n)$ for $n\notin\tilde{S}$. Then $\tilde{b}=\tilde{\chi}_{\tilde{S}}^{\tilde{c}\tilde{a}}$, and so $b\in A_{ca}$, but clearly $b\neq c$. ∎
@@
Next, fix $a,b$ and $c$ in ${\mathcal{M}}$ with $(a,b)$ and $(a,c)$ in ${\mathcal{Z}}$. We claim that $\alpha^{ab}=\alpha^{ac}$. Assume otherwise and fix an ${\mathcal{I}}$-positive $S$ such that $\alpha^{ab}(S)\cap\alpha^{ac}(S)=\emptyset$. Lift $S$ to $\tilde{S}\in{\mathcal{P}}({\mathbb{N}})$. Let $d$ be constructed in such a way that:
@@
Such a $d$ exists as each graph is sufficiently random. Note that $\neg E(a,d)$. On the other hand, we have $E(d,\chi_{S}^{ab})$ and $E(d,\chi_{S}^{ac})$, hence $E(\Phi(d),\chi_{\alpha^{ab}(S)}^{\Phi(a)\Phi(b)})$ and $E(\Phi(d),\chi_{\alpha^{ac}(S)}^{\Phi(a)\Phi(c)})$. As $\alpha^{ab}(S)$ and $\alpha^{ac}(S)$ are disjoint, the union of their complements is (modulo ${\mathcal{J}}$) equal to $\mathbb{N}$. Lifting $\Phi(d),\chi_{\alpha^{ac}(S)}^{\Phi(a)\Phi(c)},\chi_{\alpha^{ab}(S)}^{\Phi(a)\Phi(b)}$ respectively, and using the fact that $E(x,y)$ if and only if $\forall^{\mathcal{J}}n\,(E(\tilde{x}(n),\tilde{y}(n)))$, we have that $E(\Phi(a),\Phi(d))$, which is a contradiction.
@@ 2.3. A general local criterion
In the following definition, it is required that the set ${\mathcal{Z}}$ is first-order definable and that there is a first-order definition of a system of Boolean algebras that is correctly interpreted in every reduced product of models of $T$. While the theory $T^{*}$ has other models, we do not put any requirements on the interpretation of these first-order definitions in models of $T^{*}$ that are not reduced products.
@@
Let $T$ be a first order theory, and suppose that ${\mathcal{Z}}$ is a set of pairs definable in $T^{*}$. We say that $T$ recognizes coordinates at ${\mathcal{Z}}$ if for all $(a,b)\in{\mathcal{Z}}$ we have that $a\perp b$ and the sets $A_{ab}$ (as in (2.6)), as well as their canonical Boolean algebra structure, are uniformly definable131313‘Uniformly definable’ means that a single formula with parameters $a$ and $b$ defines $A_{ab}$ for all relevant pairs $a,b$ in all relevant reduced products, and that analogous fact applies to Boolean operations in the canonical Boolean algebra structure on $A_{ab}$. in all reduced products of models of $T$ over atomless ideals.
@@
Let $T$ be a theory that recognizes coordinates at a nonempty definable set ${\mathcal{Z}}$. Let ${\mathcal{M}}_{n}$ and ${\mathcal{N}}_{n}$ be models of $T$, and let ${\mathcal{I}}$ and ${\mathcal{J}}$ be atomless ideals on ${\mathbb{N}}$. Let ${\mathcal{M}}=\prod_{n}{\mathcal{M}}_{n}/{\mathcal{I}}$ and ${\mathcal{N}}=\prod_{n}{\mathcal{N}}_{n}/{\mathcal{J}}$, and suppose that $\Phi\colon{\mathcal{M}}\to{\mathcal{N}}$ is an isomorphism.